xx GRHydroInitData

GRHydroxInitxData

Luca Baiotti, Ian Hawke, Scott Hawley

24/07/2008

Abstract

GRHydroxInitxData - some initial data for GRHydro

1 Introduction This thorn generates some initial data for the GRHydro code. There are more (and more physically interesting) initial-data codes in other thorns. As with the GRHydro code itself, please feel free to add, alter or extend any part of this code. However please keep the documentation up to date (even, or especially, if it’s just to say what doesn’t work).

Currently this thorn contains a few tests that should really be test suites, some shock-tube initial data, a routine to set atmopshere everywhere on the grid (for tests), a routine to read initial data from files (not very generic, tough) and a routine to set up the simple-wave initial data .

1.1 Tests There are tests of the TVD reconstruction routine and of the routines that convert between conservative and primitive variables. These all just produce output to the screen or to fort.* files. The reconstruction test outputs the function to be reconstructed and the boundary-extended values. The conservative-to-primitive test just outputs the two sets of variables. If you haven’t altered the code an if you set

eos_polytrope::eos_gammax=xxx2.0
eos_polytrope::eos_kxxxxx=x100.0

(which are the defaults), the output should be

xxxxprimitivexvariables:
xxxxrhoxxx:xxx1.29047172182043
xxxxvelxxx:xxx9.902578465178671E-004
xxxxvelyxx:xxx9.902578465178671E-004
xxxxvelzxx:xxx9.902578465178671E-004
xxxxepsxxx:xxx0.374770481293314
xxxxpressx:x166.531726481819
xxxxw_lorx:xxx1.00000147091915

The conservative to primitive to conservative test outputs the initial and final data which should agree.

1.2 Shocktube tests There are three possible shock-tube problems, referred to as Sod, Simple and Blast, with initial data

 11 Type 11  11 11    11    11 11

11 Sod 11 1 11011 1.5 110.125110110.15

11Simple111011011 20 11 1 11011
11 Blast 11 1 11011150011 1 11011

11

The shock shape can be planar (along each axis or along the main diagonal) or spherical and the position of the plane or of the center of the sphere can be chosen though parameters. If a diagonal shock is selected, the initial data is set to either the left or right state depending on where the centre of the cell falls. Cleverer routines that weight the initial data to avoid “staircasing” may be added if there is demand. For more discussion on shock tubes seex[1].

1.3 Only atmosphere For testing purposes, this routine sets all the points to the values of the atmosphere.

1.4 Simple wave This routine testes initial data for a simple wave with sinusoidal initial function for the velocity, as described in Anile, Miller, Motta, Formation and damping of relativistic strong shocks,Phys. Fluids 26, 1450 (1983).

References

[1]xxxJ.xM. Martx and E.xMxller. Numerical hydrodynamics in Special Relativity. Living Rev. Relativity, 3, 1999. [Article in online journal], cited on 31/7/01, http://www.livingreviews.org/Articles/Volume2/1999-3marti/index.html.

[2]xxxJ.xA. Font, M. Miller, W. Suen and M. Tobias. Three Dimensional Numerical General Relativistic Hydrodynamics I: Formulations, Methods, and Code Tests Phys. Rev., D61, 044011, 2000.

2 Parameters

11advectedloopxcase 10Scope: privatex01 KEYWORD

11Description: Vx=0 or not?

11 Range 10 x01 Default: Vx=0
11Vx=0  10Useful to evaluate divB deviations
11Vx/=0  10Useful to evaluate con2prim robustness in keeping Vx const.

11

11advectedloopxdela  10Scope: privatex01  KEYWORD

11Description: How to calculate Bx field from the potential Ab̂

11 Range 10 x01 Default: Exact
11Exact  10Analytic, exact closed formula applied
11Numeric  10Finite difference approximation of the derivatives applied

11

11advectedloopxtype  10Scope: privatex01  KEYWORD

11Description: 2-dimensional or 3-dimensional?

11 Range 10 x01 Default: 2D
112D  102-dimensional (Bx=0)
113D  103-dimensional (Bˆ3 =0, where Bˆ3 —— oblique cylinder axis.

11alfvenwavexpressure  10Scope: privatex01  REAL

11Description: Pxgas for the Alfven wave

11 Range 10 x01 Default: 1.0
11(0:*  10positive

11

11alfvenwavextype  10Scope: privatex01  KEYWORD

11Description: 1-dimensional or 2-dimensional?

11 Range 10 x01 Default: 1D
111D  101-dimensional
112D  102-dimensional (in x-y plane)

11

11atmospherexvel  10Scope: privatex01  REAL

11Description: Velocity of the atmosphere if non-trivial

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11attenuatexatmosphere  10Scope: privatex01  BOOLEAN

11Description: Attenuate the velocity in the atmosphere

11                         10  x01  Default: no

11

11bhxbondixposxx  10Scope: privatex01  REAL

11Description: X-coordinate of black hole in Bondi solution

11 Range 10 x01 Default: 0.0
11*:*  10anything

11

11bhxbondixposxy  10Scope: privatex01  REAL

11Description: Y-coordinate of black hole in Bondi solution

11 Range 10 x01 Default: 0.0
11*:*  10anything

11

11bhxbondixposxz  10Scope: privatex01  REAL

11Description: Z-coordinate of black hole in Bondi solution

11 Range 10 x01 Default: 0.0
11*:*  10anything

11

11bondixbetaxsonicpt  10Scope: privatex01  REAL

11Description: Plasma beta parameter at the sonic point. Calculate bondixbmag afterwards.

11 Range 10 x01 Default: 1.0
11(0:*  10positive

11

11bondixbmag  10Scope: privatex01  REAL

11Description: Bx0 parameter for magnetized Bondi

11 Range 10 x01 Default: 0.01
110:*  10Anything positive

11

11bondixbvecxmethod  10Scope: privatex01  KEYWORD

11Description: how to compute the magnetic field vector

11 Range 10 x01 Default: direct
11direct  10directly from Cartesian metric
11transform  10transform Schwarzschild solution to Kerr Schild

11bondixcentralxmass  10Scope: privatex01  REAL

11Description: Mass of central object to find Bondi solution about

11 Range 10 x01 Default: 1.0
11(0:*  10positive

11

11bondixcentralxspin  10Scope: privatex01  REAL

11Description: Dimensionless spin of central object within Bondi solution

11 Range 10 x01 Default: 0.0
11(-1.:1.)  10dimensionless spin so any real number between -1 and 1

11

11bondixcoordinates  10Scope: privatex01  KEYWORD

11Description: Which coordinate system to use

11 Range 10 x01 Default: Isotropic
11Boyer-Lindquist  10Schwarzschild or Boyer-Lindquist (Cartesian) Coordinates
11Kerr-Schild  10Kerr-Schild (Cartesian) Coordinates
11Isotropic  10Isotropic (Cartesian) Coordinates

11

11bondixevolvexonlyxannulus 10Scope: privatex01  BOOLEAN

11Description: reset to initial data outside of bondixfreezexinnerxradius and bondixfreezexouterxradius

11                         10  x01  Default: no

11

11bondixfreezexinnerxradius 10Scope: privatex01  REAL

11Description: reset to initial at radii below this

11 Range 10 x01 Default: -1.
11*:*  10any value

11

11bondixfreezexouterxradius 10Scope: privatex01  REAL

11Description: reset to initial at radii above this

11 Range 10 x01 Default: 1e300
11*:*  10any value

11

11bondixoverwritexboundary 10Scope: privatex01  BOOLEAN

11Description: reset data to initial data in outer boundary in boundary condition

11                         10  x01  Default: no

11

11bondixradialxoffset  10Scope: privatex01  REAL

11Description: redefine rxgrid=rxKS-r0 to avoid singularity on grid

11 Range 10 x01 Default: 0.0
110:*  10Any positive number

11

11bondixrmax  10Scope: privatex01  REAL

11Description: Largest radius in units of central mass at which the solution is found

11 Range 10 x01 Default: 400.
11(0:*  10dimensionless outer radius for Bondi solution

11

11bondixrmin  10Scope: privatex01  REAL

11Description: Smallest radius in units of central mass at which the solution is found

11 Range 10 x01 Default: 1.e-15
11(0:*  10dimensionless inner radius for Bondi solution

11bxxinit  10Scope: privatex01  REAL

11Description: Initial B-field in the x-dir

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11byxinit  10Scope: privatex01  REAL

11Description: Initial B-field in the y-dir

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11bzxinit  10Scope: privatex01  REAL

11Description: Initial B-field in the z-dir

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11changexshockxdirection  10Scope: privatex01  BOOLEAN

11Description: Change the shock direction

11                         10  x01  Default: no

11

11cylxpressxinner  10Scope: privatex01  REAL

11Description: pressure in inner core

11 Range 10 x01 Default: 1.d0
11(0:*  10any positive number

11

11cylxpressxouter  10Scope: privatex01  REAL

11Description: pressure in outer region

11 Range 10 x01 Default: 3.d-5
11(0:*  10any positive number

11

11cylxrxinner  10Scope: privatex01  REAL

11Description: Inner Radius

11 Range 10 x01 Default: 0.8
11(0:*  10Any positive number

11

11cylxrxouter  10Scope: privatex01  REAL

11Description: Outer Radius

11 Range 10 x01 Default: 1.0
11(0:*  10Any positive number

11

11cylxrhoxinner  10Scope: privatex01  REAL

11Description: density in inner core

11 Range 10 x01 Default: 1.d-2
11(0:*  10any positive number

11

11cylxrhoxouter  10Scope: privatex01  REAL

11Description: density in outer region

11 Range 10 x01 Default: 1.d-4
11(0:*  10any positive number

11densxinit  10Scope: privatex01  REAL

11Description: Initial conserved mass density

11 Range 10 x01 Default: 1.29047362
11(0:*  10Anything positive.

11

11epsxinit  10Scope: privatex01  REAL

11Description: Initial specific internal energy

11 Range 10 x01 Default: 1.0d-6
11(0:*  10Anything positive.

11

11gxxxinit  10Scope: privatex01  REAL

11Description: Initial xx metric componenent

11 Range 10 x01 Default: 1.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11gxyxinit  10Scope: privatex01  REAL

11Description: Initial xy metric componenent

11 Range 10 x01 Default: 0.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11gxzxinit  10Scope: privatex01  REAL

11Description: Initial xz metric componenent

11 Range 10 x01 Default: 0.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11gyyxinit  10Scope: privatex01  REAL

11Description: Initial yy metric componenent

11 Range 10 x01 Default: 1.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11gyzxinit  10Scope: privatex01  REAL

11Description: Initial yz metric componenent

11 Range 10 x01 Default: 0.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11gzzxinit  10Scope: privatex01  REAL

11Description: Initial zz metric componenent

11 Range 10 x01 Default: 1.0
11*:*  10Anything, but be carefull to set a positive definite 3-metric!

11

11mdotxsonicptxbondi  10Scope: privatex01  REAL

11Description: Accretion rate at sonic point in hydro units

11 Range 10 x01 Default: 12.566370614359172954
11(0:*  10positive

11

11monopolexpointxbx  10Scope: privatex01  REAL

11Description: Pointlike Monopole Bx value

11 Range 10 x01 Default: 1.0
11*:*  10Any number

11monopolextype  10Scope: privatex01  KEYWORD

11Description: Which kind of monopole?

11 Range 10 x01 Default: Point
11Point  10Single point with Bx /= 0
11Gauss  10Gaussian w/radius RxGauss
111dalt  101-d alternating
112dalt  102-d alternating
113dalt  103-d alternating

11

11nxbondixpts  10Scope: privatex01  REAL

11Description: Number of points to use in determining global Bondi solution

11 Range 10 x01 Default: 2000
11(1:*  10number of points in global Bondi solution

11

11numxbondixsols  10Scope: privatex01  INT

11Description: Number of central masses about which to calculate Bondi solutions

11 Range 10 x01 Default: 1
111:100  10positive

11

11poloidalxaxb  10Scope: privatex01  REAL

11Description: Vector potential strength

11 Range 10 x01 Default: 0.1
11*:*  10Anything.

11

11poloidalxnxp  10Scope: privatex01  INT

11Description: Vector potential strength

11 Range 10 x01 Default: 3
110:*  10Any positive integer.

11

11poloidalxpxcut  10Scope: privatex01  REAL

11Description: Pressure used to confine the B field inside a star

11 Range 10 x01 Default: 1.0e-8
11(0:*  10Anything positive.

11

11poloidalxpxp  10Scope: privatex01  INT

11Description: Index of pressure factor

11 Range 10 x01 Default: 1
11(0:*  10Any non-negative integer

11

11poloidalxrhoxmax  10Scope: privatex01  REAL

11Description: Maximum initial density

11 Range 10 x01 Default: 1.0e-3
11(0:*  10Anything positive.

11

11pressxinit  10Scope: privatex01  REAL

11Description: Initial pressure

11 Range 10 x01 Default: 6.666666666666667d-7
11(0:*  10Anything positive.

11rxgauss  10Scope: privatex01  REAL

11Description: Radius for a Gaussian monopole

11 Range 10 x01 Default: 1.0
110:*  10Any positive number

11

11rxsonicptxbondi  10Scope: privatex01  REAL

11Description: Radial distance of the sonic point from the black hole in units of massxbhxbondi

11 Range 10 x01 Default: 8.0
11(0:*  10positive

11

11rhoxinit  10Scope: privatex01  REAL

11Description: Initial rest mass density

11 Range 10 x01 Default: 1.0d-6
11(0:*  10Anything positive.

11

11rotorxbvcxl  10Scope: privatex01  REAL

11Description: intial component of Bvec[0]

11 Range 10 x01 Default: 1.0
11*:*  10any real number

11

11rotorxbvcyl  10Scope: privatex01  REAL

11Description: intial component of Bvec[1]

11 Range 10 x01 Default: 0.0
11*:*  10any real number

11

11rotorxbvczl  10Scope: privatex01  REAL

11Description: intial component of Bvec[2]

11 Range 10 x01 Default: 0.0
11*:*  10any real number

11

11rotorxpressin  10Scope: privatex01  REAL

11Description: initial pressure inside rotor

11 Range 10 x01 Default: 1.d0
11(0:*  10any positive number

11

11rotorxpressout  10Scope: privatex01  REAL

11Description: initial pressure outside rotor

11 Range 10 x01 Default: 1.d0
11(0:*  10any positive number

11

11rotorxrxrot  10Scope: privatex01  REAL

11Description: radius of rotor

11 Range 10 x01 Default: 0.1
11(0:*  10any positive number

11

11rotorxrhoin  10Scope: privatex01  REAL

11Description: initial density inside rotor

11 Range 10 x01 Default: 10.d0
11(0:*  10any positive number

11rotorxrhoout  10Scope: privatex01  REAL

11Description: initial density outside rotor

11 Range 10 x01 Default: 1.d0
11(0:*  10any positive number

11

11rotorxrsmoothxrel  10Scope: privatex01  REAL

11Description: Define the radius in relative terms if so

11 Range 10 x01 Default: 0.05
11(0:*  10any positive number

11

11rotorxusexsmoothing  10Scope: privatex01  BOOLEAN

11Description: Smooth the edge?

11                         10  x01  Default: yes

11

11rotorxvxmax  10Scope: privatex01  REAL

11Description: Maximum velocity

11 Range 10 x01 Default: 0.995
11(-1:1)  10any subluminal speed (negative is clockwise)

11

11rotorxxc  10Scope: privatex01  REAL

11Description: center of rotation

11 Range 10 x01 Default: 0.5
11*:*  10Any location

11

11rotorxyc  10Scope: privatex01  REAL

11Description: center of rotation

11 Range 10 x01 Default: 0.5
11*:*  10Any location

11

11setxbondixbetaxsonicpt  10Scope: privatex01  BOOLEAN

11Description: Set plasma beta parameter instead of bondixbmag

11                         10  x01  Default: no

11

11shockxcase  10Scope: privatex01  KEYWORD

11Description: Simple, Sod’s problem or other?

11 Range 10 x01 Default: Sod
11Simple  10GRAstroxHydro test case
11Sod  10Sod’s problem
11Blast  10Strong blast wave
11Balsaralike1  10Hydro version of Balsara Test #1
11Balsara0  10Balsara Test #1, but unmagnetized
11Balsara1  10Balsara Test #1
11Balsara2  10Balsara Test #2
11Balsara3  10Balsara Test #3
11Balsara4  10Balsara Test #4
11Balsara5  10Balsara Test #5
11Alfven  10Generical Alfven Test
11Komissarov1  10Komissarov Test #1
11Komissarov2  10Komissarov Test #2
11Komissarov3  10Komissarov Test #3
11Komissarov4  10Komissarov Test #4
11Komissarov5  10Komissarov Test #5
11Komissarov6  10Komissarov Test #6
11Komissarov7  10Komissarov Test #7
11Komissarov8  10Komissarov Test #8
11Komissarov9  10Komissarov Test #9

11shockxradius  10Scope: privatex01  REAL

11Description: Radius of sperical shock

11 Range 10 x01 Default: 1.0
110.0:*  10Anything positive

11

11shockxxpos  10Scope: privatex01  REAL

11Description: Position of shock plane: x

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11shockxypos  10Scope: privatex01  REAL

11Description: Position of shock plane: y

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11shockxzpos  10Scope: privatex01  REAL

11Description: Position of shock plane: z

11 Range 10 x01 Default: 0.0
11*:*  10Anything

11

11shocktubextype  10Scope: privatex01  KEYWORD

11Description: Diagonal or parallel shock?

11 Range 10 x01 Default: xshock
11diagshock  10Diagonal across all axes
11diagshock2d  10Diagonal across x-y axes
11xshock  10Parallel to x axis
11yshock  10Parallel to y axis
11zshock  10Parallel to z axis
11sphere  10spherically symmetric shock

11

11simplexwavexconstantxcx0 10Scope: privatex01  REAL

11Description: The cx0 constant in Anile Miller Motta, Phys.Fluids. 26, 1450 (1983)

11 Range 10 x01 Default: 0.3
110:1  10It is the sound speed where the fluid velocity is zero

11

11simplexwavexvxmax  10Scope: privatex01  REAL

11Description: The vxmax constant in Anile Miller Motta, Phys.Fluids. 26, 1450 (1983)

11 Range 10 x01 Default: 0.7
110:1  10It is the maximum velocity in the initial configuration (see p. 1457, bottom of first column)

11

11sxxinit  10Scope: privatex01  REAL

11Description: Initial x component of conserved momentum density

11 Range 10 x01 Default: 0.166666658
11*:*  10Anything.

11

11syxinit  10Scope: privatex01  REAL

11Description: Initial y component of conserved momentum density

11 Range 10 x01 Default: 0.166666658
11*:*  10Anything.

11szxinit  10Scope: privatex01  REAL

11Description: Initial z component of conserved momentum density

11 Range 10 x01 Default: 0.166666658
11*:*  10Anything.

11

11tauxinit  10Scope: privatex01  REAL

11Description: Initial conserved total energy density

11 Range 10 x01 Default: 0.484123939
11(0:*  10Anything positive.

11

11usexc2pxwithxentropyxeqn 10Scope: privatex01  BOOLEAN

11Description: Use the con2prim routine that uses the entropy equation instead of the energy equation

11                         10  x01  Default: no

11

11velxxinit  10Scope: privatex01  REAL

11Description: Initial x velocity

11 Range 10 x01 Default: 1.0d-1
11*:*  10Anything.

11

11velyxinit  10Scope: privatex01  REAL

11Description: Initial y velocity

11 Range 10 x01 Default: 1.0d-1
11*:*  10Anything.

11

11velzxinit  10Scope: privatex01  REAL

11Description: Initial z velocity

11 Range 10 x01 Default: 1.0d-1
11*:*  10Anything.

11

11bvecxevolutionxmethod  10Scope: shared from HYDROBASEx01KEYWORD

11

11entropyxevolutionxmethod 10Scope: shared from HYDROBASEx01KEYWORD

11

11initialxavec  10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges:

11poloidalmagfield  10Poloidal Magnetic Field

11

11initialxbvec  10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges:

11shocktube  10Shocktube type
11cylexp  10Poloidal Magnetic Field
11poloidalmagfield  10Poloidal Magnetic Field
11magnetized Bondi  10radial magnetic field appropriate for Bondi test

11

11initialxentropy  10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges:

11magnetized Bondi  10Initial entropy for a radial magnetic field appropriate for Bondi test

11initialxhydro  10Scope: shared from HYDROBASEx01  KEYWORD

11Extends ranges:

11shocktube  10Shocktube type
11shocktubexhot  10Shocktube with hot nuclear EOS
11onlyxatmo  10Set only a low atmosphere
11readxconformal  10After reading in initial alp, rho and gxx from h5 files, sets the other quantities
11simplexwave  10Set initial data from Anile Miller Motta, Phys.Fluids. 26, 1450 (1983)
11monopole  10Monopole at the center
11cylexp  10Cylindrical Explosion
11rotor  10”Magnetic Rotor test from DelZanna,Bucciantini , and Londrillo A&A 400, 397-413 (2003)”
11advectedloop  10Magnetic advected loop test
11alfvenwave  10Circularly polarized Alfven wave
11see [1] below  10Spherical single black hole Bondi solution
11see [1] below  10Spherical single black hole Bondi solution - TEST ISO CASE!!!!!!
11see [1] below  10Magnetized Spherical single black hole Bondi solution
11see [1] below  10Magnetized Spherical single black hole Bondi solution - TEST ISO CASE!!!!!!

11

hydro\_bondi\_solution

[1]

hydro\_bondi\_solution\_iso

[1]

magnetized\_bondi\_solution

[1]

magnetized\_bondi\_solution\_iso

11temperaturexevolutionxmethod 10Scope: shared from HYDROBASEx01KEYWORD

11

11timelevels  10Scope: shared from HYDROBASEx01INT

11

11yxexevolutionxmethod  10Scope: shared from HYDROBASEx01KEYWORD

11

3 Interfaces

General Implements:

grhydroxinitxdata

Inherits:

grhydro grid eosxomni

Grid Variables

3.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x x01 Details x11x
11grhydroxinitxdataxreflevel10GRHydroxinitxdataxreflevelx01  compact110
11                      10  x01 description11Refinement level GRHydro is working on right now
11                      10  x01 dimensions110
11                      10  x01 distribution11CONSTANT
11                      10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                      10  x01 timelevels111
11                      10  x01variable type11INT

11

3.0.2 PROTECTED GROUPS

11 x Group Names x 10x Variable Names xx01 Details x11x
11simplexwavexgridxfunctions10  x01  compact110
11                        10  simplextmp  x01 description111D arrays for the simple-wave routine
11                        10  cxs  x01 dimensions113
11                        10  x01 distribution11DEFAULT
11                        10  x01 group type11GF
11 10 x01 tags11checkpoint=”no”
11                        10  x01 timelevels111
11                        10  x01variable type11REAL
11 simplexwavexscalars 10  x01  compact110
11                        10  simplexrhox0  x01 description11values at v=0
11                        10  simplexepsx0  x01 dimensions110
11                        10  x01 distribution11CONSTANT
11                        10  x01 group type11SCALAR
11                        10  x01 timelevels111
11                        10  x01variable type11REAL
11 simplexwavexoutput 10  x01  compact110
11                        10  simplexrho  x01 description11output variables for the simple-wave routine
11                        10  simplexeps  x01 dimensions113
11                        10  x01 distribution11DEFAULT
11                        10  x01 group type11GF
11 10 x01 tags11checkpoint=”no”
11                        10  x01 timelevels111
11                        10  x01variable type11REAL

11

Uses header:

SpaceMask.h

4 Schedule This section lists all the variables which are assigned storage by thorn EinsteinEvolve/GRHydroxInitData. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11x11Conditional:
11x 11GRHydroxinitxdataxreflevel
11x 11GRHydroxinitxdataxreflevel
11x 11GRHydroxinitxdataxreflevel
11x 11GRHydroxinitxdataxreflevel
11x 11simplexwavexgridxfunctions
11x 11simplexwavexscalars
11x 11simplexwavexoutput

11x 11x

11

Scheduled Functions

CCTKxPARAMCHECK x grhydroxinitdataxcheckparameters x check parameters x

11x11Language:11c
11 x 11Type: 11function

11

HydroBasexInitial (conditional)

x grhydroxmonopolem x monopole initial data x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional)

x grhydroxshocktube x shocktube initial data x

11x11Language:11fortran
11 x 11Type: 11function

HydroBasexInitial (conditional) x grhydroxcylindricalexplosionm x cylindrical explosion initial data - mhd-only x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxinitxdataxrefinementlevel x calculate current refinement level x

11x11Before: 11grhydroxcon2primtest
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxcon2primtest x testing the conservative to primitive solver x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxinitxdataxrefinementlevel x calculate current refinement level x

11x11Before: 11c2p2cxcall
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x c2p2cm x testing conservative to primitive to conservative - mhd version x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x c2p2c x testing conservative to primitive to conservative

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxinitxdataxrefinementlevel x calculate current refinement level x

11x11Before: 11p2c2pxcall
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x p2c2pm x testing primitive to conservative to primitive - mhd version x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x p2c2p x testing primitive to conservative to primitive x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxrotorm x mhd rotor initial data x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxinitxdataxrefinementlevel x calculate current refinement level x

11x11Before: 11p2c2pxcall
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x p2c2pmxpolytype

x testing primitive to conservative to primitive - mhd polytype version

x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxreconstructionxtest x testing the reconstruction x

11x11Language:11fortran
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Storage: 11grhydroxprimxbext
11 x 11x  11grhydroxscalars
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxonlyxatmo x only atmosphere as initial data x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxreadconformaldata x set the missing quantities, after reading in from file initial data from

conformally-flat codes (garching) x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxsimplewave x set initial data from anile miller motta, phys.fluids. 26, 1450

(1983)

x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x grhydroxsimplewavexanalysis x compute some output variables for the simple wave

11x11After: 11grhydroxentropy
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxbondixiso x setup grhydro vars for the hydrodynamic bondi solution x

11x11After: 11hydrobasexexcisionmasksetup
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxbondimxiso x setup grhydro vars for the magnetized bondi solution x

11x11After: 11hydrobasexexcisionmasksetup
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxbondi x setup grhydro vars for the hydrodynamic bondi solution x

11x11After: 11hydrobasexexcisionmasksetup
11 x 11Language:11c
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxadvectedloopm x mhd advected loop initial data x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxbondim x setup grhydro vars for the magnetized bondi solution x

11x11After: 11hydrobasexexcisionmasksetup
11 x 11Language:11c
11 x 11Type: 11function

HydroBasexCon2Prim (conditional)

x grhydroxbondimxrange x force analytic solution outside anulus x

11x11Before: 11con2prim
11 x 11Language:11c
11 x 11Type: 11function

11

HydroBasexBoundaries (conditional)

x grhydroxbondimxboundary x force analytic solution in boundaries x

11x11Before: 11hydrobasexselectxboundaries
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxINITIAL (conditional)

x grhydroxpoloidalmagfieldm x set up a poloidal magnetic field. it expects the other fluid variables

already to be set, as for example in the tov solution x

11x11After: 11hydrobasexinitial
11 x 11Before: 11grhydrotransformprimtolocalbasis
11 x 11Language:11fortran
11 x 11Type: 11function

11

CCTKxINITIAL (conditional)

x hydrobasexboundaries x call boundary conditions after magnetic field initial data setup x

11x11After: 11grhydroxpoloidalmagfieldm
11 x 11Before:11grhydrotransformprimtolocalbasis
11 x 11Type: 11group

11

HydroBasexInitial (conditional)

x grhydroxalfvenwavem x circularly polarized alfven wave initial data x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional)

x grhydroxshocktubexhot x hot shocktube initial data

11x11After: 11hydrobasexyxexone
11 x 11x  11hydrobasexzero
11 x 11Language:11fortran
11 x 11Type: 11function

11

HydroBasexInitial (conditional) x grhydroxshocktubem x shocktube initial data - mhd version x

11x11Language:11fortran
11 x 11Type: 11function

11

ApplyBCs (conditional)

x grhydroxdiagshockxboundarym x diagonal shock boundary conditions x

11x11After: 11boundaryconditions
11 x 11x  11boundary::boundaryxclearselection
11 x 11Language:11fortran
11 x 11Sync:  11grhydro::dens
11 x 11x  11grhydro::tau
11 x 11x  11grhydro::scon
11 x 11x  11grhydro::bcons
11 x 11x  11grhydro::psidc
11 x 11Type: 11function

11

ApplyBCs (conditional)

x grhydroxdiagshockxboundarym x diagonal shock boundary conditions x

11x11After: 11boundaryconditions
11 x 11x  11boundary::boundaryxclearselection
11 x 11Language:11fortran
11 x 11Sync:  11grhydro::dens
11 x 11x  11grhydro::tau
11 x 11x  11grhydro::scon
11 x 11x  11grhydro::bcons
11 x 11Type: 11function

11

ApplyBCs (conditional)

x grhydroxdiagshock2dxboundarym x 2-d diagonal shock boundary conditions x

11x11After: 11boundaryconditions
11 x 11x  11boundary::boundaryxclearselection
11 x 11Language:11fortran
11 x 11Type: 11function

Aliased Functions x

11Alias Name: xxxxxxx11Function Name:
11c2p2c  11c2p2cxcall
11c2p2cM  11c2p2cxcall
11p2c2p  11p2c2pxcall
11p2c2pM  11p2c2pxcall
11p2c2pMxpolytype  11p2c2pxcall